1.5, 2.5, 2.5, 4.1, 5, 6.7, 8.12, 9.9, 10.5, 10.4



Determine the mean and standard deviation

1
Expert's answer
2021-12-09T13:44:09-0500

n= 10

"\mu=\frac{\sum}{n}"

"\mu=\frac{1.5+2.5+.....+10.4}{10}"


"\mu =6.122"

Std dev

"\sqrt{\sum_i^n\frac{(x_i-\bar x^2)}{n-1}}"

"\sqrt{\sum\frac{(1.5-6.122)^2+.....(10.4-6.122)^2)}{9}}"

=3.4804